Ruth has decided to drop her collision insurance because her car is getting old. Her total annual premium is $916, of which $170.60 covers collision insurance
1
Expert's answer
2020-05-18T17:15:21-0400

After Ruth drops her collusion insurance, her annual premium will be:

"\text{Premium}=960-170.6"


"\text{Annual Premium}= 745.4"


The premium that paid quarterly are:

"\text{Quaterly payments}=\dfrac{\text{Annual premiums}}{4}"


"\text{Quaterly payments}=\dfrac{745.6}{4}=186.4"